Journalist Report – February 28th

Sol 16 – An extraordinary routine

“They began to plan people’s lives and libraries; they began to instruct and push about the very people who had come to Mars to get away from being instructed and ruled and pushed about.”

– Chapter 16 of The Martian Chronicles by Ray Bradbury

When we awoke and found that the Martian desert was once again covered in snow, and having heard the winds howl in the middle of the night, our first reflex was to reach for the binoculars, strategically placed on the main windowsill, from which we could see the atmospheric instruments. We breathed: rocks, tensors, and metal rods hammered into the ground had kept everything upright! There was no need for an emergency EVA.

This morning, we decided to change it up a bit: instead of our usual workout session prepared by Corentin, we relaxed and stretched to the sound of a yoga session brought by Jérémy. Feeling a little bit too relaxed and sleepy afterwards, we put on some music, did some push-ups and just goofed around the Hab. It’s these sparse moments of unbridled fun and laughs that convince me our crew works: no amount of hard work or accumulated stress can prevent us from having a good time all together. In the past few days, the song “Cheerleader” by OMI started coming back more and more often during our morning activities; I think it might end up being Crew 275’s mission song!

In the Upper Deck of the Hab, I watched this morning as crewmembers came and went, to and from different parts of the station, for their cognitive tests, experiments, routine activities, in the Science Dome or Observatory, in pairs or alone. We have definitely settled in a routine, one that we will only realize was extraordinary once we return to Earth!

After eating a delicious focaccia prepared with dill and tomatoes from the GreenHab, Adrien, Quentin, and I went to the Science Dome to prepare our second EVA to Candor Chasma. We had one hour to study the 3D map of the canyon and try to visualize the location of all 10 checkpoints, before we go out on EVA tomorrow. I was surprised to see how differently I worked the problem compared to when I only had the 2D map: because we could “navigate” inside the canyon and see different views of rock formations, Adrien and I started sketching the location of the checkpoints in the order we would find them when exploring the canyon.

By late afternoon, the winds were not too high and the sky was clear enough to open the dome of the solar observatory, and Alexandre was able to observe our Sun for the first time! The telescope is equipped with a Hydrogen-alpha filter, making it possible to look at the Sun’s chromosphere without damaging the eyes. It also allows to see certain phenomena, such as prominences, sunspots, and flares on the solar surface.

Back at the Hab, I saw Quentin working on solving some issues with the location tracking system. Half the crew is now “tagged”, the other half still pending… I am also proud to announce that Crew 275 has the best possible Health and Safety Officer: to help me feel less restless in the evenings, he prepared an entire workout session just for me!

Astronomy Report – February 28th

Astronomy Report

Name: Alexandre VINAS

Crew: 275
Date: 02/28/2023


Robotic Telescope Requested (choose one) MDRS-14 MDRS-WF

Objects Viewed: /

Problems Encountered: /


Solar Features Observed: Sunspot

Problems Encountered: Two attempts to observe the Sun. The first one, the sky was too cloudy. The second one, at the sunset, I could visually observe the Sun but I did not have time to take any pictures.

GreenHab Report – February 28th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 70.2°F

Average humidity: 22%

Hours of supplemental light: 0 hour

Daily water usage for crops: 13.21 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 148.73 gallons

Time(s) of watering for crops: 9:38 & 17:25

Changes to crops: N/A

Narrative: Yesterday I used what I harvested to put some fresh salad into our fajitas and prepare a good focaccia with the dill and tomatoes.

We will use today’s fresh spinach harvest to cook the famous ‘cream spinach’ recipe you can find in the MDRS cookbook!

In the GreenHab today less work was required, but what a pleasure it is to water and maintain crops every day. The bean I planted few days ago is already popping out, and it feels great to see that.

There wasn’t that much sun today, so it wasn’t hot in there, but plants still needed to be well watered.

My aquaponics system is doing really well, apart from the parsley that doesn’t really like it, I can clearly notice how faster than usual plants grow.

Harvest: 161 g of spinach, 1.138 kg of tomatoes, 7 g of basil

Support/supplies needed: N/A

Operations Report – February 28th

Crew 275 Operations Report 28Feb2023
SOL: 16

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• GreenHab Carbon Monoxide alarm

Notes on non-nominal systems: None

Spirit rover used: No
• Hours: 209.6
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.5
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.4
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations.
• WATER USE: 11.33
• Water: 446.49 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS. MDRScampus is used for experimental purpose.

Summary of suits and radios: Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations.
• WATER USE: 13.21 gallons
• Heater: On
• Supplemental light: No
• Harvest: 161 g of spinach – 1138 g of tomatoes – 7 g of basil

Summary of ScienceDome operations: The next EVA for the photogrammetry experiment was prepared by the EVA Crew.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM today.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support:
• It is a bit cold for working comfortably in the Science Dome, when do you think the Dual Split could be turned back on?
• The kitchen knobs that were delivered last week do not seem to work because they are too large to fit in the dedicated rods.

Sol Summary Report – February 28th

Crew 275 Sol Summary Report 28Feb2023

Sol: 16

Summary Title: An extraordinary routine

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

With a probability of precipitation of 0%, I must say that we were all surprised to see that the surroundings were covered with snow when we woke up. Since no EVA was planned for today, we took some time to relax a bit in the morning with a longer sport session including 30 minutes of yoga.

With the anchors of the location tracking system fully deployed, we have started equipping the crew with tags that they have to always keep on their arm or on their belt. This system enables us to check who is in which part of the station at which time. Later on, it will allow the scientists to correlate these locations and the environmental conditions at these locations with other parameters such as stress levels extracted from heart rate variability.

With the yesterday’s harvest, everyone is delighted to have the opportunity to eat fresh food. Our Crew Botanist and our Health & Safety Officer have prepared some delicious meals that recharged our batteries. We then discussed about more profound topics, which we did not really have the opportunity to do so far. These small things are extremely important for the mood, as we have already entered our third week here, and the end of the mission still looks relatively far.

In the afternoon, the crewmembers assigned to tomorrow’s EVA have taken some time to prepare for their mission with a 3D render of Candor Chasma. Tomorrow, they will only be allowed to read their notes to find the checkpoints defined during the previous EVA there.

Our Crew Astronomer also took the opportunity of today’s weather to visually observe the Sun. Unfortunately, he could not record any video of his observation. On his first attempts, the Sun hid behind some clouds, while on the second attempt, it hid behind the landscape, as it was already getting late.

Look Ahead Plan:

We plan an EVA at Candor Chasma for tomorrow (Sol 17) and another on Friday (Sol 19). The EVA crews will have to find specific checkpoints based on a 2D map or a 3D render. We are also planning to have a second geology EVA on Sol 20, to further test the MetMet instrument. We would like to take the opportunity of these different EVAs to change the batteries of the weather instruments, since they cannot last for more than two days, especially in cold weather.

The BBC is planned to arrive at the MDRS on Friday (Sol 19) between 8:30 and 9:00. They would follow us on our EVA to Candor Chasma, then film inside the station in the afternoon, and leave at around 17:00.

Anomalies in work: None.

Weather: Windy, low -5°C/23°F, high 7°C/44°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Request, Operations Report, Journalist

Report, GreenHab Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site